OpenStudy (anonymous):

2yx + 10y=z and it turns into (z/2y)-5= x can anyone explain me the steps??

jhonyy9 (jhonyy9):

2yx +10y = z factoriz out 2y on left side 2y(x+5) =z divide both sides by 2y x+5 = z/(2y) subtract 5 from both sides x = (z/2y) - 5 hope so much that is understandably sure right easy good luck bye

OpenStudy (anonymous):

in the secomd step, you divided 10y by 2yx?